# Memory-aware feedback enhances power in active information engines

### v1

@@ I Introduction
In this work, we develop a stochastic thermodynamic framework for an information engine operating in an active bath with temporally correlated noise. We propose a proportional adjustment feedback protocol that partially retains the system’s memory, allowing feedback to better align with the underlying dynamics. Using a harmonic trap under overdamped Langevin dynamics, we show that this memory-aware control strategy enhances work extraction compared to conventional feedback protocol. Our results reveal that temporal correlations can be harnessed to improve engine performance. This study extends the scope of information thermodynamics to active and temporally structured environments and highlights the utility of adaptive, memory-aware feedback in nonequilibrium control.
@@ II Model
where the overdot denotes time derivative. The parameters $\gamma$, $\kappa$ and $\lambda$ represent the damping constant, the stiffness of the harmonic trap, and the center of potential, respectively. All variables with subscript $t$ evolve in time. The thermal noise satisfies $\langle\xi_{t}\rangle=0$, $\langle\xi_{t}\xi_{t^{\prime}}\rangle=2\gamma k_{{}_{\text{B}}}T\delta(t-t^{\prime})$ where $k_{{}_{\text{B}}}$ is the Boltzmann constant and $T$ is the temperature of the thermal bath, consistent with the Einstein relation. The active noise satisfies $\langle\eta_{t}\rangle=0$, $\langle\eta_{t}\eta_{t^{\prime}}\rangle=A^{2}\exp(-|t^{\prime}-t|/\tau_{c})$, where $A$ and $\tau_{c}$ are the parameters representing the amplitude and correlation time of the noise, respectively [25, 23]. The thermal and active noises are statistically independent, as ensured by $\langle\xi_{t}\eta_{t^{\prime}}\rangle=0$.
Figure 1: Operation of an information engine and memory effect in active baths. (a) Along the time line, measurement and feedback occur at $t=n\tau_{-}$ and $t=n\tau_{+}$, respectively. In the left panel (Measurement), the inverted triangle indicates the measured position $\hat{x}_{t}$ of the particle in the absolute frame, while the arrow ($\hat{y}_{t}$) represents the same measurement position in the relative frame. In the right panel (Feedback), the center of the potential is instantaneously shifted from $\lambda_{t_{-}}$ to $\lambda_{t_{+}}$ as part of the feedback process. (b) The left and right panels display contour plots of the joint probability distribution between the relative position $y$ and the active noise $\eta$ at $n\tau_{-}$ and $n\tau_{+}$, respectively. The distributions are evaluated in the periodic steady state. (c) The heatmap shows a joint probability distribution $p(y_{t},\eta_{t})$ obtained from numerical simulations using active Ornstein-Uhlenbeck (AOUP) process. The black dots and the connecting line represents the conditional average of the active noise given the position $y$.
@@
If the measurement is free from error, the protocol $\lambda_{t_{+}}=\hat{x}_{t}$ (i.e., $\hat{y}_{t}=0$) clearly yields the most exploitative update, and this has been used as a practical experimental ansatz [16]. On the other hand, when the measurement error is present, the Bayesian estimate $\hat{x}_{t}$ offers a more refined update compared to the direct use of raw measurement [29]. In active baths, feedback based on error-free measurements has been assumed to be the most exploitative protocol [22, 30]. However, this assumption breaks down as such feedback disrupts memory in active baths.
Quantifying memory in a meaningful way requires consideration of both the system and its intended function [31, 32]. In our case, we are interested in the statistics of $W_{t}$, whose average is determined by $\langle x_{t}^{2}\rangle$, where $\langle\cdot\rangle$ denotes an ensemble average. During the relaxation process, the statistics of $x_{t}$ is coupled with those of the active noise $\eta_{t}$. However, this correlation is erased immediately after the feedback, as illustrated in Fig. 1(b).
@@
Finally, we propose a proportional-shift feedback scheme to mitigate memory loss during the feedback process. A similar protocol has previously been studied in a different context, that is, to test the thermodynamic uncertainty relation in thermal baths [36]. In the present work, we investigate how modulating the feedback influences its interplay with the system’s relaxation dynamics, and how this, in turn, impacts work extraction, as illustrated in Fig. 2(a). The feedback is implemented as
@@
where $g$ is the proportional gain parameter, with the standard full-shift feedback (FSF) being recovered when $g=1$.
@@ III Control of relaxation dynamics: Memory-assisted relaxation
Figure 2: Proportional adjustment as a protocol for memory preservation. (a) The inverted triangle indicates the measured position of the Brownian particle at time $t$, and the solid arrow represents a feedback with (attenuated) gain $g$. (b) Time evolution of the covariance $C_{t}$ between the relative position $y_{t}$ and the active noise $\eta_{t}$: blue for $g=1$ and magenta for $g=0.5$. The initial covariance is taken from the steady state of the cyclic operation, and increases as the proportionality $g$ deviates from 1. The black dashed line denotes the asymptotic covariance in the limit $\tau\to\infty$.
@@
Importantly, we emphasize that the power and average work of the system depend solely on the covariance and do not require higher-order moment contributions. Thus, the covariance serves as a complete and sufficient metric to quantify the relationship between memory retention and the power output of an information engine.
@@ IV Optimal performance metrics in active baths
Figure 3: Profiles of average work extraction per cycle and its classification under varying control parameters. (a) Heatmap of the average work extracted per cycle as a function of the cycle period $\tau$ and feedback gain $g$. The black solid line in the right panel marks the zero-work boundary. The orange curve on each panel indicates the optimal feedback gain gg* for each value of $\tau$. As the cycle period shortens, gg* decreases monotonically below 1, deviating from the standard full-shift feedback (FSF). (b) Average extracted work per cycle as a function of $\tau$ for different combinations of feedback gain and measurement error: cyan, blue, olive-green, and magenta curves correspond to $(g,\sigma^{2}/V_{\xi})=(0.5,~0.5)$, (1.0,0.5)(1.0,~0.5), (1.0,1.3)(1.0,~1.3), and (1.0,2.2)(1.0,~2.2), respectively, all in identical active baths. For the olive-green curve, the black vertical marker indicates the transition point where the average work changes sign, near $\tau\sim 0.66\tau_{r}$. Dashed lines indicate the asymptotic values as $\tau\to\infty$. (c) Classification of operational regimes in the $(g,\sigma^{2})$ parameter space. The blue, magenta, and olive-green regions indicate parameter combinations where the average work is always positive, always negative, or changes sign as a function of $\tau$, respectively. The gray shaded region corresponds to settings where the engine either fails to satisfy the convergence condition of Eq. (8) or yields consistently negative work.
@@ IV.1 Work extraction efficiency under noisy feedback
However, when measurement error is present, both positive and negative work extractions may occur within individual cycles. The average extracted work per cycle is given by
$\langle W_{\text{cyc}}\rangle$ $=\frac{\kappa}{2}\left(V_{\tau_{-}}-V_{\tau_{+}}\right)$
$=\frac{\kappa}{2}g\left[2V_{\tau_{-}}-g(V_{\tau_{-}}+\sigma^{2})\right]~.$ (12)
@@
Figure 4: Profiles of power and classification under varying control parameters. (a) Power profiles are shown for different activity levels, defined by the dimensionless parameter $\alpha\equiv A^{2}/2\kappa k_{\text{B}}T$, represented by blue, orange, and green curves. Dashed lines indicate linear fits in the short-period limit ($\tau\ll\tau_{r}$), defined by the intercept $P^{(0)}$ and slope $P^{(1)}$. Each row shares a common active noise correlation time. The left column displays results for FSF, while the right column shows those for an attenuated feedback ($g=0.5$). Insets zoom into the short-period regime of the orange curves, and vertical black lines indicate the period $\tau$ at which the power is maximized. (b) All six panels include the same boundary curves, which demarcate regions where the short-period intercept $P^{(0)}$ is positive (lower left) or negative (upper right). Yellow curves indicate the activity-dependent boundary where the short-period slope $P^{(1)}$ changes sign, positive in the green/white regions and negative in the blue regions.
@@ V Discussion and Conclusion
Our findings underscore the importance of feedback-relaxation interplay in the design of optimal information engines. The proportional adjustment framework may be further extended to more general feedback strategies, including finite-time protocols [37]. However, such generalizations require a consistent and thermodynamically sound formulation of work and heat in active systems, where conventional definitions are no longer valid.
@@
We also examined the optimal conditions for power, power output is maximized at short (finite) cycle times. Therefore, identifying the optimal operating time is crucial for maximizing power output. These findings provide valuable insights into the development of adaptive, dynamically tunable feedback strategies for information engines, particularly for optimizing engine’s performance.
@@ Appendix B Asymptotic behavior of work and power in the fast and slow cycle limits
Figure S1: Local extrema of $\sigma_{\alpha}^{*2}(g)$. (a) Number of solutions $n(g_{\rm ext})$ to $f^{\prime}_{\alpha}(g)=0$ as a function of $\alpha$ in the regime $g\in(0,2)$. Two separate transitions occur at $\alpha=4/9$ and $\alpha=1/2$. (b) Sign of $P^{(1)}$ at $\alpha=4/9$, analogous to Fig. 4(b). The yellow curve marks $P^{(1)}=0$, and the yellow cross ($\times$) indicates its slope vanishes. (c) Location of the extremum $g_{\rm ext}$ as a function of $\alpha$. At $\alpha=4/9$, a single point bifurcates into a local minimum (blue) and a local maximum (red curve). (d) Sign of $P^{(1)}$ at $\alpha=0.49$. The yellow curve corresponds to $P^{(1)}=0$, showing both a local minimum and a local maximum. The blue (red) curve traces the trajectory of the minimum (maximum), starting from the black cross ($\times$) at $\alpha=4/9$. For $\alpha>1/2$, the minimum disappears while the maximum becomes the global maximum, as illustrated in Fig. 4(b), particularly in the panels for $\alpha=1/2$, $0.52$, and $1$.

### v2

@@ I Introduction
In this work, we develop a stochastic thermodynamic framework for an information engine operating in an active bath with temporally correlated noise. We propose a proportional adjustment feedback protocol that partially retains the system’s memory, allowing feedback to better align with the underlying dynamics. Using a harmonic trap under overdamped Langevin dynamics, we show that this memory-aware control strategy enhances work extraction compared to conventional feedback protocol. Our results reveal that temporal correlations can be harnessed to improve engine performance. This study extends the scope of information thermodynamics of active and temporally structured environments by highlighting the utility of adaptive, memory-aware feedback in nonequilibrium control.
@@ II Model
where the overdot denotes time derivative. The parameters $\gamma$, $\kappa$ and $\lambda$ represent the damping constant, the stiffness of the harmonic trap, and the center of potential, respectively. All variables with subscript $t$ evolve in time. Denoting ensemble averages by angular brackets $\langle\cdot\rangle$, thermal noise satisfies $\langle\xi_{t}\rangle=0$, $\langle\xi_{t}\xi_{t^{\prime}}\rangle=2\gamma k_{{}_{\text{B}}}T\delta(t-t^{\prime})$ where $k_{{}_{\text{B}}}$ is the Boltzmann constant and $T$ is the temperature of the thermal bath, consistent with the Einstein relation. The active noise satisfies $\langle\eta_{t}\rangle=0$, $\langle\eta_{t}\eta_{t^{\prime}}\rangle=A^{2}\exp(-|t^{\prime}-t|/\tau_{c})$, where $A$ and $\tau_{c}$ are the parameters representing the amplitude and correlation time of the noise, respectively [25, 23]. The thermal and active noises are statistically independent, as ensured by $\langle\xi_{t}\eta_{t^{\prime}}\rangle=0$.
Figure 1: (Color online) Operation of an information engine and memory effect in active baths. (a) Along the time line, measurement and feedback occur at $t=n\tau_{-}$ and $t=n\tau_{+}$, respectively. In the left panel (Measurement), the inverted triangle indicates the measured position $\hat{y}_{t_{-}}$ relative to the center of the potential. The red dashed line represents the particle density inferred from the measurement. In the right panel (Feedback), an instantaneous shift of the potential center changes the reference frame. (b) The left and right panels display contour plots of the joint probability distribution between the relative position $y$ and the active noise $\eta$ at $n\tau_{-}$ and $n\tau_{+}$, respectively. The distributions are evaluated in the periodic steady state. (c) The heatmap shows a joint probability distribution $p(y_{t},\eta_{t})$ obtained from numerical simulations using active Ornstein-Uhlenbeck (AOUP) process. The black dots and the connecting line represents the conditional average of the active noise given the position $y$.
@@
If the measurement is free from error, the protocol $\lambda_{t_{+}}=\hat{x}_{t}$ (i.e., $\hat{y}_{t}=0$) clearly yields the most exploitative update, and this has been used as a practical experimental ansatz [16]. On the other hand, when the measurement error is present, the Bayesian estimate of $y_{t}$ offers a more refined update compared to the direct use of raw measurement [29]. In active baths, feedback based on error-free measurements has been assumed to be the most exploitative protocol [22, 30]. However, this assumption breaks down as such feedback disrupts memory in active baths.
Quantifying memory in a meaningful way requires consideration of both the system and its intended function [31, 32]. In our case, we focus on the expectation value of $W_{t}$, determined by $\langle y_{t}^{2}\rangle$, while analysis of higher moments remains a natural and important direction, particularly in connection with fluctuation theorems [27, 22]. During the relaxation process, the statistics of $x_{t}$ is coupled with those of the active noise $\eta_{t}$. However, this correlation is erased immediately after the feedback, as illustrated in Fig. 1(b).
@@
Finally, we propose a proportional-shift feedback scheme to mitigate memory loss during the feedback process. As illustrated in Fig. 2(a), the feedback is implemented as
@@
where $g$ is the proportional gain parameter, with the standard full-shift feedback (FSF) being recovered when $g=1$. Similar protocols have previously been proposed in different contexts, for example, to generate a broad family of nonequilibrium steady states via partial resetting [36], or to test thermodynamic uncertainty relations in thermal baths [37]. In the present work, we investigate how modulating the feedback influences its interplay with the system’s relaxation dynamics, and how this, in turn, impacts work extraction.
@@ III Control of relaxation dynamics: Memory-assisted relaxation
Figure 2: (Color online) Proportional adjustment as a protocol for memory preservation. (a) The inverted triangle indicates the measured position of the Brownian particle at time $t$, and the horizontal arrows denote the shift of the potential center, with magenta and blue vertical markers representing proportional adjustment and full shift, respectively. (b) Time evolution of the covariance $C_{t}$ between the relative position $y_{t}$ and the active noise $\eta_{t}$: blue for $g=1$ and magenta for $g=0.5$. The initial covariance is taken from the steady state of the cyclic operation, and increases as the proportionality $g$ deviates from 1. The black dashed line denotes the asymptotic covariance in the limit $\tau\to\infty$.
@@
In summary, by modulating the feedback adjustment gain, the engine’s cyclic steady state can be replaced with a controllable, switchable relaxation behavior. Moreover, the cyclic dynamics determine the average work extracted per cycle:
$\langle W_{\text{cyc}}\rangle=\dfrac{\kappa}{2}\left(V_{\tau_{-}}-V_{\tau_{+}}\right).$ (12)
Importantly, Eq. (12) depends solely on covariance and do not require higher-order moment contributions. Thus, covariance serves as a complete and sufficient metric to quantify the relationship between memory retention and the power output of an information engine.
@@ IV Optimal performance metrics in active baths
Figure 3: (Color online) Profiles of average work extraction per cycle and its classification under varying control parameters. (a) Heatmap of the average work extracted per cycle as a function of the cycle period $\tau$ and feedback gain $g$. The black solid line in the right panel marks the zero-work boundary. The orange curve on each panel indicates the optimal feedback gain gg* for each value of $\tau$. As the cycle period shortens, gg* decreases monotonically below 1, deviating from the standard full-shift feedback (FSF). (b) Average extracted work per cycle as a function of $\tau$ for different combinations of feedback gain and measurement error: cyan, blue, olive-green, and magenta curves correspond to $(g,\sigma^{2}/V_{\xi})=(0.5,~0.5)$, (1.0,0.5)(1.0,~0.5), (1.0,1.3)(1.0,~1.3), and (1.0,2.2)(1.0,~2.2), respectively, all in identical active baths. For the olive-green curve, the black vertical marker indicates the transition point where the average work changes sign, near $\tau\sim 0.66\tau_{r}$. Dashed lines indicate the asymptotic values as $\tau\to\infty$. (c) Classification of operational regimes in the $(g,\sigma^{2})$ parameter space. The blue, magenta, and olive-green regions indicate parameter combinations where the average work is always positive, always negative, or changes sign as a function of $\tau$, respectively. The gray shaded region corresponds to settings where the engine either fails to satisfy the convergence condition of Eq. (8) or yields consistently negative work.
@@ IV.1 Work extraction efficiency under noisy feedback
However, when measurement error is present, both positive and negative work extractions may occur within individual cycles. Combining Eq. (11a) with Eq. (12), we obtain
$\langle W_{\text{cyc}}\rangle=\frac{\kappa}{2}g\left[2V_{\tau_{-}}-g(V_{\tau_{-}}+\sigma^{2})\right]~.$ (13)
@@
In the absence of explicit knowledge of the measurement error, the analysis can be generalized by considering its possible range. The maximum conceivable error then provides a conservative bound on the feedback gain that should not be exceeded to ensure positive work extraction. For instance, if positivity is required for all values of $\tau$, the corresponding maximal feedback gain is determined by the boundary between the olive-green and blue regions in Fig. 3(c).
Figure 4: (Color online) Profiles of power and classification under varying control parameters. (a) Power profiles are shown for different activity levels, defined by the dimensionless parameter $\alpha\equiv A^{2}/2\kappa k_{\text{B}}T$, represented by blue, orange, and green curves. Dashed lines indicate linear fits in the short-period limit ($\tau\ll\tau_{r}$), defined by the intercept $P^{(0)}$ and slope $P^{(1)}$. Each row shares a common active noise correlation time. The left column displays results for FSF, while the right column shows those for an attenuated feedback ($g=0.5$). Insets zoom into the short-period regime of the orange curves, and vertical black lines indicate the period $\tau$ at which the power is maximized. (b) All six panels include the same boundary curves, which demarcate regions where the short-period intercept $P^{(0)}$ is positive (lower left) or negative (upper right). Yellow curves indicate the activity-dependent boundary where the short-period slope $P^{(1)}$ changes sign, positive in the green/white regions and negative in the blue regions.
@@ V Discussion and Conclusion
Our findings underscore the importance of the feedback-relaxation interplay in the design of optimal information engines. The proportional adjustment framework may be further extended to more general feedback strategies, including finite-time protocols [38, 39, 40]. However, such generalizations require a consistent and thermodynamically sound formulation of work and heat in active systems, where conventional definitions are no longer valid. Despite these challenges, our findings provide guiding principles for designing resetting processes that exist under either thermal or active noise, or both [41, 42]. A related theoretical framework of partial resetting [36] may provide a complementary perspective on the efficiency of feedback-driven engines.
@@
We also examined the optimal conditions for power and found that the power output is maximized at short, finite cycle times. Hence, identifying the optimal operating time is crucial for maximizing performance. These findings offer valuable insights for developing adaptive and dynamically tunable feedback strategies in information engines, complementing recent advances in related research areas [43, 44, 45].
@@ Appendix B Asymptotic behavior of work and power in the fast and slow cycle limits
Figure S1: (Color online) Local extrema of $\sigma_{\alpha}^{*2}(g)$. (a) Number of solutions $n(g_{\rm ext})$ to $f^{\prime}_{\alpha}(g)=0$ as a function of $\alpha$ in the regime $g\in(0,2)$. Two separate transitions occur at $\alpha=4/9$ and $\alpha=1/2$. (b) Sign of $P^{(1)}$ at $\alpha=4/9$, analogous to Fig. 4(b). The yellow curve marks $P^{(1)}=0$, and the yellow cross ($\times$) indicates its slope vanishes. (c) Location of the extremum $g_{\rm ext}$ as a function of $\alpha$. At $\alpha=4/9$, a single point bifurcates into a local minimum (blue) and a local maximum (red curve). (d) Sign of $P^{(1)}$ at $\alpha=0.49$. The yellow curve corresponds to $P^{(1)}=0$, showing both a local minimum and a local maximum. The blue (red) curve traces the trajectory of the minimum (maximum), starting from the black cross ($\times$) at $\alpha=4/9$. For $\alpha>1/2$, the minimum disappears while the maximum becomes the global maximum, as illustrated in Fig. 4(b), particularly in the panels for $\alpha=1/2$, $0.52$, and $1$.